OpenStudy (anonymous):

The sum of four consecutive integers is 58. What is the largest of the four integers? A. 13 B. 16 C. 12 D. 21

OpenStudy (anonymous):

16

OpenStudy (anonymous):

add 13 plus 14 then 15 and 16 and you get 58

OpenStudy (anonymous):

thanks for the medal

OpenStudy (raden):

let that numbers are : x, x+1, x+2, x+3 the expression of algebra according information above : x+x+1+x+2+x+3=58 we simplify be 4x + 6 = 58 4x = 58-6 4x = 52 x = 52/4 = 13 so, that numbers are 13,14,15, and 16 as the larger number is 16

OpenStudy (anonymous):

or you can do what rad did